User name kuchina
Log entry time 09:44:05 on May 1,2007
Entry number 200706
Coda crashed at the beginning of the shift. ROC4 stopped responding. Remote resets wouldn't work. Ole Hansen asked for the controlles access and pushed the right button in the Hall. 9:43 beam is back to 100uA and we are starting the first run od the day shift.